OpenStudy (esmerelda):

write the equation of a line that is perpendicular to the given line that passes through the given point.

OpenStudy (esmerelda):

here is the equation y-2=7/3(x+5);(-4,9)

OpenStudy (danjs):

\[\large y-2=\frac{ 7 }{ 3 }(x+5)\] Do you know the slope of that given line?

OpenStudy (esmerelda):

no

OpenStudy (danjs):

remember the point slope form of a line.. y - y1 = m*(x - x1) the slope of that line is m, and it goes through point (x1,y1)

OpenStudy (esmerelda):

A. y-9=-3/7(x-4) B. y-9=3/7(x-4) C. y-9=-3/7(x+4) D. y-9=3/7(x+4)

OpenStudy (esmerelda):

these are the answer choices

OpenStudy (danjs):

The given line has the slope 7/3. To get a slope perpendicular to that, you flip over the fraction and change the sign.

OpenStudy (danjs):

A perpendicular line will have a slope m=-3/7 and goes through given point (x1,y1)=(-4 , 9)

OpenStudy (danjs):

Point slope form of a line... y - y1 = m*(x - x1) sub in the values

OpenStudy (esmerelda):

is it A

OpenStudy (danjs):

almost, for the (x-x1) part though, it is minus a negative 4, turns into a plus 4. y - 9 = -3/7 (x - (-4)) y - 9 = -3/7 (x + 4)

OpenStudy (esmerelda):

oh thanks